### v4

@@ Abstract
Associations between phenotype and genomic and epigenomic markers are often derived by correlation. Systems Biology aims to make more robust connections and uncover broader insights by modeling the cellular mechanisms that produce a phenotype. The question of choosing the modeling methodology is of central importance. A model that does not capture biological reality closely enough will not explain the system’s behavior. At the same time, highly detailed models suffer from computational limitations and are likely to overfit the data. Boolean networks strike a balance between complexity and descriptiveness and thus have received increasing interest. We previously described an algorithm for fitting Boolean networks to high-throughout experimental data that finds the optimal network with respect to the information in a given dataset. In this work, we describe a simple extension that enables the modeling of asynchronous dynamics, i.e. different reaction times for different network nodes. In addition, we present a new method for pseudo-time assignment for single-cell RNA sequencing data that is derived from the modeling procedure. Our approach greatly simplifies the construction of Boolean network models for time-series datasets, where asynchronicity often occurs. We demonstrate our methodology by integrating real data from transcriptomics experiments. These results significantly expand the applicability of the Boolean network model to experimental data.
@@
Boolean Network, Inference, Gene Expression, Biological Network1 Introduction
Cellular processes that alter the functional profile of the cell are subject to complex, adaptable regulation. This regulation is achieved by networks of interactions between molecules that integrate environmental signals and the current state of the cell. In order to understand these networks, various modeling methodologies have been proposed, varying in their degree of complexity and expressiveness(Karlebach and Shamir, 2008). Due to limited knowledge of the reactions involved in carrying out each regulatory interaction, Boolean models, first introduced by Kauffman, are frequently chosen as a general-purpose modeling methodology. Several questions arise when reconstructing a Boolean network from data. First, as with any computational task, one must ask whether the process can be accomplished efficiently (Karlebach and Shamir, 2012). The second question is how to find the correct trade-off between model size and the fit to the data. Finally, a method to map continuous or discrete measurements into binary values is needed (Hofensitz201). In addressing the first question, Karlebach and Shamir showed that the problem of optimally fitting a Boolean model to binary data is NP-Complete (Karlebach and Shamir, 2008), making a solution impractical in the worst case and likely challenging for many other instances. The second question is usually addressed by independently minimizing the fit of individual regulatory logic tables to the binarized values of their targets. Bayesian approaches have also been proposed, but these arguably deviate from the simplicity of the Boolean model by introducing a probabilistic parameters. Karlebach and Robinson proposed a non-probabilistic criterion that is based on Kolmogorov Complexity (Kolmogorov, 1968) ,and takes into account both the fit to the data and the total size of the model (Karlebach and Robinson, 2023a), and proposed an algorithm for optimizing according to it. As for the third question, while a gold-standard has not been agreed upon, various methodologies have been developed that generate useful mappings in practice (Berestovsky and Nakhleh, 2013). An often overlooked challenge in modeling is accounting for variation in reaction times. A gene may be activated at a different rate than another gene that has a regulatory association to it, and this can result in a time-series dataset where not all regulations take effect at every sampling point. A related issue occurs when consecutive sampling times capture the same network state repeatedly. Since the data is usually affected by noise, it may be hard to know when the exact same state has been sampled twice. To address these issues, our goal in this paper is to extend the methodology described in (Karlebach and Robinson, 2023a) for Boolean networks with asynchronous dynamics, thereby obtaining a rigorous reconstruction methodology that can handle time-series data where the sampling rate has not been optimized, or where the rates of different regulatory interactions differ significantly. In the following section, we specify and extend our methodology and explain its application to asynchronous network modeling. We also show how heuristics that address the computational complexity of the problem can be used in the case of asynchronous dynamics. Finally, we discuss pseudo-time assignment to single-cell RNA-Sequencing experiments that is based on the modeling methodology. In the Results section we describe modeling of real data with asynchronous Boolean networks. In the Conclusion section we summarize our findings and assess the impact of the method on future modeling efforts.
2 Data and Methods
In this section we will assume that the network reconstruction uses transcriptomics data, although the methodology is applicable to any high-throughput dataset. Additionally, all the values are assumed to have been binarized, i.e. mapped to the set of Boolean values 0 and 1. Mapping of real data to Boolean values is demonstrated in the Results section. A Boolean network can be described by a directed graph G(V,E), where G is a set of nodes or genes, and E are edges such that a gene ’s regulators are the nodes from which directed edges extend to , i.e. . Each set of regulators of a gene is associated with a logic function that has Boolean inputs and one Boolean output. The Boolean inputs are determined when the nodes that are associated with the function’s inputs are assigned Boolean values. Biologically, the input nodes correspond to regulators, and their combined activity or inactivity determines the Boolean value of their target gene. Given a Boolean assignment to all the nodes in the network, also referred to as a state, the next state can be computed by combining the outputs of all the logic functions. A sequence of states that are derived from one another in this way are called a trajectory. If a state is always followed by an identical state in a trajectory, it is called a steady state. In experimental data, a steady state is usually provided as one measurement per gene or network node. The dynamics of the network are synchronous if the values of all the genes are updated by the logic functions simultaneously. If updates can be delayed for any subset of the genes, for any number of network state traversals, we will say the the dynamics are asynchronous. In gene expression datasets, the logic is often unknown and regulators are at best known approximately. A gene expression dataset consists of a N M matrix, where N corresponds to the number of genes whose expression level was measured, and M corresponds to the number of experiments. The entry at indices i,j contains a Boolean value that is equal to 1 if the gene is in the active state, and otherwise equal to 0. If the data consists of steady states, every row of the matrix represents a network state in which none of the genes will change their activity without external perturbation. If the data consists of trajectories, sequences of rows correspond to consecutive time points in which the network updates the activity states of some or all of its genes. The data can also consist of both steady states and trajectories. In addition to the gene expression dataset, the modeler determines a set of plausible regulatory interactions. For example, the modeler can identify those genes whose protein products bind to the promoter of a target gene, and refine the selection by intersecting this information with correlations in the expression dataset. From this initial set, the modeler want to choose the optimal one, including the logic tables by which the regulators determine the state of the target. An expression dataset will, in practice, always contain some noise. This means that a single set of regulators taking the same value in two network states will have a different effect on the target in each of these states. An inference methodology must therefore identify where such inconsistencies are a result of experimental noise, and where they are the result of asynchronous dynamics. The approach described in the next section minimizes the sum of such inconsistencies and the number of bits in the network encoding. To briefly state the rationale, every additional network bit doubles the number of possible networks, and similarly every mismatch halves the number of remaining solutions to fit. Hence, a non-random solution will result in a sum that is significantly smaller than the size of the dataset (Karlebach and Robinson, 2023b).
@@ 2 Data and Methods
In order to find an optimal solution to the problem, we formulate it as 0/1 Integer programming. First, we assume that the number of regulators of each gene is relatively small, in the sense that the logic table defining the regulation is of manageable size. The variables of the problem are denoted by uppercase English letters. A variable is defined for every measurement, i.e. an entry in the expression matrix that describes a gene and its activity at a given state, and is equal to 1 if there is a mismatch between the observed value of the gene at that measurement and the value that the model assigns it. An variable is defined for every combination of regulator values, and is equal to 1 if the state of the target gene is set to 1 for that combination, and otherwise it is equal to 0. An variable is defined for every potential regular-target pair. It is equal to 1 if the regulator is chosen for the optimal model, and otherwise to 0. A variable is defined for every gene and possible number of regulators for that gene, and is equal to 1 if the gene has at least that number of regulators, and otherwise it is equal to 0. At the end of the section we will also define a variable, which will allow us to implement asynchronous dynamics. Using these variables, we first describe the constraints of the model, and then the objective function: Let denote the observed Boolean value of gene at experiment . The corresponding variable is , and it is equal 1 if the value of gene in experiment does not match the model’s assignment, and otherwise 0. If is the index of a steady state in the data, is a gene with regulators , we go over every possible combination of values for these regulators , and for each one add the following constraint:
@@
Next, for every gene and each one of its regulators , we create a Boolean variable . In other words, every potential regulator of gene is associated with an variable for that gene. For every two different assignment of values to ’s regulators, i.e. inputs to the logic function that sets the value of , the sum of variables of regulators which have different values in the two assignments is constrained to be greater than the differences between the variables that determine the outputs for these assignments. For example, with two regulators and and two assignments and to the variables respectively, must be greater than and also greater than . If the outputs for these two assignments are different, only the change in can explain this difference,as has the same value in both assignments. More generally, this constraint means that two different outputs can never occur for the exact combination of regulatory inputs, for otherwise the regulatory logic is not a function. The variable, which is defined for gene and every possible number of regulators of that gene, is constrained to be greater than if or if . Now we can set the weights of variables in the objective to match the inference criterion: First, a weight of 1 is given to variables. Now if regulators are chosen for a gene, all its variables will be set to 1. Therefore, we set the weight of the first variable of the gene to be the log base 2 of the number of ways to choose a first regulator plus the log base 2 of the number of logic tables possible for one regulator. We then set the weight of the second variable of the gene to be the log base 2 of the number of ways to choose a second regulator after the first one was already chosen plus the log base 2 of the number of logic tables with two regulators, minus the log base 2 of the number of logic table with one regulator. So the cost of encoding the logic tables cancel out by consecutive variables, while the cost of choosing the regulators is produced by the combination of all the that are set to 1. If we denote the number of logic tables with regulators as , the weight of the variable is set to So far, we assumed that the updates of the model are synchronous. We now adapt the 0/1 IP formulation to fit asynchronous dynamics, by modifying how it models trajectories. We add a new type of variable called the variable. This variable is defined for every constraint that involves the variables in a trajectory, as defined in (1) and (2). It is added to the right hand side of the constraint, and therefore if it is equal to 1 it allows the output of the logic function to not agree with its inputs. We further constrain the variable to be smaller than 1 minus the differences between the chosen value of target gene at the state at which the regulatory effect is taking effect and the previous state, i.e., the value selected for the gene by the model at these states. This constraint only allows the output of the logic function to disagree with its input if the output does not change, i.e. if the regulatory update is not immediate. Using the same notation as before, the additional constraints on the variable can be described as follows:
@@
We set the weight of every variable to 1 in the objective function. Consider a network that is the optimal solution for some dataset . If it sets the value of the variable to 1 at some time , then the corresponding target must exert a regulatory change on one of its own targets after the delay introduced by the variable, for otherwise a better solution could have been obtained without setting the variable. Therefore, the trajectory of the model when there is no delay (i.e. when the variable is not set) is different than the one it uses in the optimal solution. Now if we set the variable to 0 instead, set the suffix of the trajectory from time to fit exactly the trajectory of the model that was fit to from time and afterwards, then must also be optimal for this new trajectory . If not, and there is another better fit model, then when flipping the bits back, it will still be better than on the original dataset, whether uses the variable or not. Therefore, like the variables, every variable is equivalent to one bit in the encoding of the network. As is the case for any 0/1 Integer Programming formulation, the value of the objective function is a sum of the weights of variables that are set to 1 in the solution. Finally, for each target gene we constrain the first variable in each trajectory to be smaller or equal to the sum of the target’s variables, such that the target would only be able to use the variables if it has regulators assigned to it. Powerful solvers like Gurobi have dramatically improved our ability to solve 0/1 Integer Programming problems. However, custom heuristics offer even better performance than those devised for the general problem. We now describe such heuristics.
@@
Perhaps the simplest heuristic for a trajectory is to perform a single pass over the data, state by state starting from the first state, and to record every input-output pair observed as long as it does not conflict with pairs observed before it. When a conflict happens, the value of the target gene is flipped to match the output that was previously observed. A more sophisticated approach was suggested by Karlebach and Robinson (Karlebach and Robinson, 2023a), and can be applied to an expression data sets composed of either steady states or equal-length trajectories:
1.Choose a set of regulators.
2.If the set has a single steady state, return it as a solution.
@@
3.If the set has s single trajectory, solve any inconsistencies using the single-pass heuristic, and return it as a solution.
@@
4.If the size of the set of steady states or trajectories is larger than 1 but the set is consistent with the regulators, return that set of states as a solution, possibly removing some redundant regulators by backward elimination.
@@
5.Otherwise, cluster the states and round the cluster centers into Boolean vectors, then solve the problem recursively for the cluster centers. This generates a set of consistent states S. For every state in the original set, choose its closest neighbor in S, and flip its values one by one to match the neighbor’s values until all inconsistencies with states in S have been resolved, or until it is equal to the neighbor, which is already consistent. At that point add it to S so it can be compared to states that have not been made consistent yet. At the end of the process, return S excluding the cluster centers.
@@
The set of regulators in step 1 can be chosen from the current LP solution, for example all regulators which have a value of at least 0.5. The number of clusters should be such that the clusters contain either one state or more than two, since a cluster center for two Boolean states may contain multiple values of 0.5 whose rounding to 0 or 1 is arbitrary. If the dataset contains both steady states and trajectories, then the recursive heuristic can be run for the steady states, and then the resulting logic can be used to remove inconsistencies from the trajectory using the single-pass heuristic. If trajectories have different lengths, equal-sized contiguous subsequences of trajectories can be solved by the recursive heuristic, and the remaining inconsistencies then resolved by the single-pass heuristic. Care should be taken that clustering of these subsequences is biologically meaningful, for otherwise poor solutions may be result due to their incompatibility.
@@
To complete the modeling methodology, the heuristic needs to be adapted for asynchronous dynamics. If a gene’s value does not match the output expected by the values of its regulators, but it is consistent with the value of the gene in the previous time point, then it is no longer flagged as an inconsistency. Additionally, when fixing inconsistencies by performing a pass over trajectories and building a set of logic functions, functions are only updated when their target genes change their values between consecutive time steps. With these changes, the heuristic can apply to asynchronous trajectories, or a combination of steady states and such trajectories.
We have integrated this heuristic into our 0/1 Integer Programming implementation in order to improve the upper bound on the nascent solution.

### v5

@@ Abstract
The network inference problem arises in biological research when one needs to quantitatively choose the best protein-interaction model for explaining a phenotype. The diverse nature of the data and nonlinear dynamics pose significant challenges in the search for the best methodology. In addition to balancing fit and model complexity, computational efficiency must be considered. In this paper, we present a novel approach that finds a solution that fits the observed dataset and otherwise a minimal number of unobserved datasets. We present algorithms for computing Boolean networks that optimally satisfy this criterion, and allow for asynchronicity network dynamics. Furthermore, we show that using our methodology a solution to the pseudo-time inference problem, which is pertinent to the analysis of single-cell data, can be intertwined with network inference. Results are described for real and simulated datasets.
@@
Boolean Network, Inference, Optimization, Time-series1 Introduction
Research in molecular biology often aims to reveal a mechanistic understanding of underlying processes, in contrast to correlations between observations. In particular, one aims to describe cellular networks of interactions between proteins that generate a specific phenotype [1, 2, 3, 4]. Often, these networks display complex, nonlinear dynamics due to combinatorial interactions between regulators and feedback loops in the topology. [5]. A key aspect of the problem is the Curse of Dimensionality, which ensures that the number of networks and datasets grows exponentially with the number of involved proteins, making inference susceptible to overfitting.
Gene regulatory networks, which control the expression of genes in the cell, are an important type of biological networks that are studied using transcriptomic technologies. These technologies measure the level (expression) of all the genes in a sample and, more recently, in single cells [6], allowing the detection of regulatory relationships between them. When the measurement is at the sample level, we refer to the data as RNA-Sequencing data (RNA-Seq), and when it is at the single-cell level as single-cell RNA-Seq (scRNA-Seq).
Numerous, fundamentally different inference methodologies have been proposed. For example, Keyl et al. used an explainable AI approach and layer-wise relevance propagation in order to select gene regulators that are ranked as important predictors in a neural network model [7]. SCODE is a method for inferring ordinary differential equations from scRNA-Seq data, such that the regulatory relationships can be derived from the equations [8]. CEFCON uses a graph neural network with an attention mechanism for constructing the regulatory network from prior network knowledge, scRNA-Seq data and trajectory information [9]. Chen et al. used a genetic algorithm to reconstructed a Probabilistic Boolean Network from scRNA-Seq data, a model that accommodates probabilistic transitions in a Boolean network model [10]. However, none of the existing methodologies satisfy all of the following criteria: define an objective criterion for comparing any pair of networks, find optimal solution with respect to this inference criterion, regularize any solution as part of the inference criterion and process both steady-state and time-series data. Presenting such a methodology is the goal of this paper.
In fitting a model that can be described using discrete quantities, and in the absence of dataset noise, the best solution is the one that uses the shortest encoding, since with the strictest limit on encoding length one can fit the smallest number of unobserved datasets. This minimal length encoding is known as the Kolmogorov Complexity [11]. In the case of biological datasets, noise is practically always present, but the same principle can be used for inference in the presence of noise. We will show that this leads to an inference methodology that satisfies the requirements described above, using Boolean networks as the modeling language [12, 13, 14].
Another challenge that arises specifically when working with scRNA-Seq is that individual measurements originate from a time-series but a temporal order is not provided with the dataset. In other words, we obtain "snapshots" of the cell at different states, that may have occurred consecutively in time and thus describe a network trajectory, but we do not know at which point in the trajectory each state, or snapshot, occurred. Here, too, many efforts have been devoted to the development of inference methods such that the cell order, known as pseudo-time, could be derived. To name a few examples, Riba et al. developed DeepCycle, a method that predicts position relative to the cell cycle based on the relationship between spliced and unspliced transcripts. DeepCycle trains an autoencoder that encodes the relationship between spliced and unspliced scRNA-Seq reads as the transcriptional phase of the cell, a latent variable that maps a cell into a position on the periodic cell cycle trajectory [15]. Liu et al. introduced reCAT, a method for cell-cycle stage inference that finds a cycle that minimizes transcriptional differences between adjacent cells [16]. The CONFESS method utilizes image information to reconstruct single-cell dynamics [17]. PAGA creates a map connecting cell groups based on the estimation of the connectivity of manifold partitions [18]. GraphDDP combines clusters into differentiation trajectories using a forced-based graph layout approach [19]. We argue that the pseudo-time inference problem is closely related to the network inference problem and that a network model should be utilized to solve it. We will demonstrate how this can be done in the context of the novel methodology suggested here.
2 Methods
2.1 Network Inference
Discussing the inference of a Boolean network model from gene expression data requires some basic terminology. A state (of a biological cell) corresponds to a multidimensional vector, where each vector entry is the level of one gene. This level describe the degree to which a gene is active, and for our purposes it can take only two values - 0 (inactive) and 1 (active). A Boolean network can be viewed as a function that maps a Boolean state to another, potentially identical, state. For a given network, a sequence of states such that each state maps to its successive state in the sequence is called a trajectory. A trajectory whose states are repeated indefinitely by the network is known as an attractor. An attractor consisting of a single state is referred to as a steady state. By definition, steady states and trajectories correspond to a specific model. To provide some biological context, changes in gene expression during cellular differentiation will result in a trajectory that can be modeled using a gene regulatory network, described as a Boolean network. The trajectories and steady states generated by a network are known as the network dynamics.
A dataset is a collection of measurements of trajectories and steady states. These measurements can be affected by noise, in which case states do not perfectly match the mapping defined by the network because some gene levels are incorrectly measured (noise bits). The goal of the inference problem is to find the network that generated the dataset. It should be clarified here that there are no hidden variables in the problem - the value of every gene is always observed. This is a property of the experimental techniques that generate the data. While adding unobserved variables does not overly complicate the methodology, it is not required for our purposes.
A network can be described concisely by a directed graph G(V,E), where G is a set of nodes or genes, and E are edges such that a gene ’s regulators are the nodes from which directed edges extend to , i.e. . Each set of  regulators of a gene is associated with a logic function that has  Boolean inputs and one Boolean output. The Boolean inputs are determined when the nodes that are associated with the function’s inputs are assigned Boolean values. Biologically, the input nodes correspond to regulators, and their combined activity or inactivity determines the Boolean value of their target gene. Given a Boolean assignment to all the nodes in the network, its state is defined, and the next state can be computed by combining the outputs of all the logic functions.
If the change in gene levels can be deferred, such that it is not immediately applied to the consecutive state, we say that the network is asynchronous. In that case, the function defined by the network as a whole is non-deterministic. Otherwise, the network is said to be synchronous.
A gene expression dataset as a N  M matrix, where N corresponds to the number of genes whose expression level was measured, and M corresponds to the number of experiments. The entry at indices i,j contains a Boolean value that is equal to 1 if the gene is in the active state, and otherwise equal to 0. Initially, we assume that the order between states is known for every trajectory. This is the case for biological datasets in which the cells have been synchronized or in which the cells are at steady state. An algorithm for inferring pseudo-time follows from the basic methodology, and will be detailed later on.
We will also assume that a set of plausible regulatory interactions has been determined for each gene. This can be achieved using other experimental technologies like ChIP-Seq [20], or by keeping regulators whose levels correlate with their target’s.
From this initial set, the modeler want to choose the optimal subset, including the logic tables by which the regulators determine the state of the target. For mapping biological measurements to Boolean values we will use existing methods [21, 22, 23, 24], and in this section we assume that all gene values are already Boolean.
Were the data noiseless, we would seek a solution whose binary encoding is as short as possible, since the number of networks that can be described by such an encoding is minimal, and hence the solution will fit a minimal amount of unobserved data [11]. Every bit doubles the number of networks that the encoding can describe. In practice, datasets are always affected by noise. We therefore need a criterion for combining noise bits and network bits such that the solution will match a minimal number of unobserved datasets. Karlebach and Robinson [25] showed, in the context of synchronous dynamics, that minimizing the sum of network encoding bits and noise bits is equivalent to minimizing the number of unobserved datasets that can be matched. A generalization to asynchronous dynamics will be discussed shortly.
In order to find an optimal solution to the problem, we formulate it as 0/1 Integer programming.
@@ 2.1 Network Inference
The variables of the problem are denoted by uppercase English letters. A  variable is defined for every measurement, i.e. an entry in the expression matrix that describes a gene and its activity at a given state, and is equal to 1 if there is a mismatch between the observed value of the gene at that measurement and the value that the model assigns it. An  variable is defined for every combination of regulator values, and is equal to 1 if the state of the target gene is set to 1 for that combination, and otherwise it is equal to 0. An  variable is defined for every (potential-regulator,target) pair. It is equal to 1 if the regulator is chosen in the optimal solution, and otherwise to 0. A  variable is defined for every gene and possible number of regulators for that gene, and is equal to 1 if the gene has at least that number of regulators, and otherwise it is equal to 0. Later in this section we will also define a  variable, which will allow us to implement asynchronous dynamics.
Using these variables, we first describe the constraints of the model, and then the objective function:
Let  denote the observed Boolean value of gene  at experiment . The corresponding  variable is , and it is equal to 1 if the value of gene  in experiment  does not match the model’s assignment, and otherwise 0. If  is the index of a steady state in the data,  is a gene with regulators , we go over every possible combination of values for these regulators  ,  and for each combination add the following constraint:
@@
Next, for every gene and each one of its regulators , we create a Boolean variable . In other words, every potential regulator of gene is associated with an variable for that gene. For every two different assignments of values to ’s regulators, i.e. inputs to the logic function that sets the value of , the sum of variables of regulators which have different values in the two assignments is constrained to be greater than the differences between the variables that determine the outputs for these assignments. For example, with two regulators and and two assignments and to the variables, respectively, must be greater than and also greater than . If the outputs for these two assignments are different, only the change in can explain this difference,as has the same value in both assignments. More generally, this constraint means that two different outputs can never occur for the exact combination of regulatory inputs, for otherwise the regulatory logic is not a function. The variable, which is defined for gene and every possible number of regulators of that gene, is constrained to be greater than the mean of the gene’s R variables minus if or if , where is the number of candidate regulators (or variables) of . To set the weights of variables in the objective to match the inference criterion, a weight of 1 is given to variables. Now if regulators are chosen for a gene, all its variables will be set to 1. Therefore, we set the weight of the first variable of the gene to be the logarithm (base 2) of the number of ways to choose a first regulator plus the of the number of logic tables possible for one regulator. We then set the weight of the second variable of the gene to be the of the number of ways to choose a second regulator after the first one was already chosen plus the of the number of logic tables with two regulators, minus the of the number of logic tables with one regulator. So the costs of encoding the logic tables cancel out by consecutive variables, while the cost of choosing the regulators is produced by the combination of all the that are set to 1. If we denote the number of logic tables with regulators as , the weight of the variable is set to
So far, we assumed that the updates of the model are synchronous. We now adapt the 0/1 IP formulation to fit asynchronous dynamics. We add a new type of variable called the variable. This variable is defined for every constraint that involves the variables in a trajectory, as defined in (1) and (2). It is added to the right hand side of the constraint, and therefore if it is equal to 1 it allows the output of the logic function to not agree with its inputs. We further constrain the variable to be smaller than 1 minus the differences between the chosen value of target gene at the state at which the regulatory effect is taking effect and the previous state, i.e., the values selected for the gene by the model at these states. The latter constraint only allows the output of the logic function to disagree with its input if the output does not change, i.e. if the regulatory update is not immediate. Using the same notation as before, the additional constraints on the variable can be described as follows:
Finally, for each target gene we constrain the first  variable in each trajectory to be smaller or equal to the sum of the target’s  variables, such that the target would only be able to use the  variables if it has at least one regulator assigned to it.
@@
We set the weight of every  variable to 1 in the objective function. Consider a network  that is the optimal solution for some dataset . If it sets the value of the  variable to 1 at some time , then the corresponding target must exert a regulatory change on one of its own targets after the delay introduced by the  variable, for otherwise a better solution could have been obtained without setting the  variable. Therefore, the trajectory of the model when there is no delay (i.e. when the  variable is not set) is different than the one it uses in the optimal solution.
Now if we set the  variable to 0 instead, set the suffix of the trajectory from time  to fit exactly the trajectory of the model that was fit to  from time  and afterwards, then  must also be optimal for this new trajectory . If not, and there is another better fit model, then when flipping the bits back, it will still be better than  on the original dataset, whether  uses the  variable or not. Therefore, like the  variables, every  variable is equivalent to one bit in the encoding of the network. The value of the objective function is a sum of the weights of variables that are set to 1 in the solution.
Powerful solvers like Gurobi [26] have dramatically improved our ability to solve 0/1 Integer Programming problems. Custom heuristics can be integrated with the solver to improve performance. We now describe such heuristics.
@@
Perhaps the simplest heuristic for a trajectory is to perform a single pass over the data, state by state starting from the first state, and to record every input-output pair observed as long as it does not conflict with pairs observed before it. When a conflict occurs, the value of the target gene is flipped to match the output that was previously observed. A more sophisticated approach was suggested by Karlebach and Robinson [25], and can be applied to an expression data set composed of either steady states or equal-length trajectories:
1. Choose a set of regulators.
2. If the set has a single steady state, return it as a solution.
@@
3. If the set has s single trajectory, solve any inconsistencies using the single-pass heuristic, and return it as a solution.
@@
4. If the size of the set of steady states or trajectories is larger than 1 but the set is consistent with the regulators, return that set of states as a solution, possibly removing some redundant regulators by backward elimination.
@@
5. Otherwise, cluster the states and round the cluster centers into Boolean vectors, then solve the problem recursively for the cluster centers. The recursive call returns a set of consistent states S. For every state in the original set, choose its closest neighbor in S, and flip its values one by one to match the neighbor’s values until all inconsistencies with states in S have been resolved, or until it is equal to the neighbor, which is already consistent. At that point add it to S so it can be compared to states that have not been made consistent yet. At the end of the process, return S excluding the cluster centers.
@@
The set of regulators in step 1 can be chosen from the current LP solution, for example all regulators which correspond to an  variable with value of at least 0.5. If the dataset contains both steady states and trajectories, then the recursive heuristic can be run for the steady states, and then the resulting logic can be used to remove inconsistencies from the trajectories using the single-pass heuristic. If trajectories have different lengths, equal-sized contiguous subsequences of trajectories can be solved by the recursive heuristic, and the remaining inconsistencies then resolved by the single-pass heuristic. Care should be taken that clustering of these subsequences is biologically meaningful, for otherwise poor solutions may be result due to their incompatibility.
@@
It remains to adapt the heuristic to allow for asynchronous dynamics. In the adapted version, if a gene’s value does not match the output expected by the values of its regulators, but it is consistent with the value of the gene in the previous time point, then it is no longer flagged as an inconsistency. Additionally, when fixing inconsistencies by performing a pass over trajectories and building a set of logic functions, functions are only updated when their target genes change their values between consecutive time steps. With these changes, the heuristic can be applied to asynchronous trajectories, or a combination of steady states and such trajectories.
As a final note, the heuristic described in 1 is in fact a family of heuristics. The order in which Boolean values are flipped in step 4 does not have to be arbitrary, and different criteria for this order will result in different solutions. Similarly, there are multiple ways to select a set of regulators, and various ways to cluster the states. Different choices will result in different heuristics.
2.2 Pseudo-time